You found a treasure map, and it's said: staring at well walk 30 m north, 20 m 30° N of East, 15 m 18° south of west, and finally 10 m north. How far is the treasure map from the well?


1
Expert's answer
2021-09-30T18:27:53-0400

"x=20\cdot\cos30°-15\cdot\cos18°=3.055\ (m)"


"y=30+20\\cdot\\sin30\u00ad\u00b0-15\\cdot\\sin18\u00b0+10=45.365\\ (m)"


"L=\sqrt{3.055^2+45.365^2}\approx45.5\ (m)" . Answer